# FastDDS-Based Middleware System for Remote X-Ray Image Classification Using Raspberry Pi

### v1

@@ Abstract
Index Terms:
IoT, Fast-DDS, ResNet50, Real-time Systems, DDS, X-Ray, Healthcare.
publicationid: pubid: 0000–0000/00\$00.00 © 2024 IEEE
I Introduction
@@
The rest of the paper is as follows. Section II reviews the approaches regarding middleware technologies, including state-of-the-art vision-based deep learning models diagnosing chest diseases. Section III explains the proposed methodology, while Section IV illustrates the experimental setup. Section V shows the results and discussions, while Section VI discusses the conclusion and future work.
II Literature Review
@@ V Results and Discussions
VI Conclusion and Future Work
@@
Acknowledge
The authors would like to acknowledge all support provided by King Fahd University of Petroleum & Minerals (KFUPM).

### v2

# Article Type: Feature Article

@@ Abstract
journal: Publication Name
[…]
@@
The rest of the paper is as follows. Section LITERATURE REVIEW reviews the approaches regarding middleware technologies, including state-of-the-art vision-based deep learning models diagnosing chest diseases. Section METHODOLOGY explains the proposed methodology, while Section EXPERIMENTAL SETUP illustrates the experimental setup. Section RESULTS AND DISCUSSIONS shows the results and discussions, while Section CONCLUSION AND FUTURE WORK discusses the conclusion and future work.
LITERATURE REVIEW
@@
CONCLUSION AND FUTURE WORK
@@
ACKNOWLEDGMENTS
The authors would like to acknowledge all support provided by Alfozan Academy, the Center for Excellence in Development of Nonprofit Organizations (CEDNPO), and King Fahd University of Petroleum and Minerals (KFUPM).